# Temperature is All You Need for Generalization in Langevin Dynamics and other Markov Processes

### v1

@@ 3 Special Case: Continuous Langevin Dynamics
Clearly, given Theorem 2.7 all we need to do in order to derive explicit generalization bounds for any Markovian training procedure, is to find a stationary distribution, and bound its potential (or its expectation at $p_{0}$). In this section, we will exemplify our results in a few special cases of continuous-time Langevin dynamics (CLD), which have a normalized stationary distribution, and we can write it analytically.
@@
where $L_{S}\geq 0$ is twice continuously differentiable, $\mathbf{w}_{t}$ is a standard Brownian motion, and $\mathbf{r}_{t}$ is a reflection process that constrains ${\boldsymbol{\theta}}_{t}$ within $\Theta$. Such weight clipping is quite common in practical scenarios such as NN training. For simplicity, we assume that $\mathbf{r}_{t}$ has normal reflection, meaning that the reflection is perpendicular to the boundary. An established result in the analysis of SDERs states that under these assumptions (9) has a unique stationary distribution $p_{\infty}\left({\boldsymbol{\theta}}\right)\propto e^{-\beta L_{S}\left({\boldsymbol{\theta}}\right)}\mathbb{I}_{\Theta}\left\{{\boldsymbol{\theta}}\right\}$ (see Section H.2). Thus, when $p_{0}=\mathrm{Uniform}\left(\Theta\right)$, we have $p_{0}=\nu$.
@@
    $\mathbb{E}_{{\boldsymbol{\theta}}_{t}\sim p_{t}}\left[E_{D}({\boldsymbol{\theta}}_{t})-E_{S}({\boldsymbol{\theta}}_{t})\right]\leq\sqrt{\frac{\beta\mathbb{E}_{p_{0}}L_{S}({\boldsymbol{\theta}})+\ln\left(1/\delta\right)}{2N}}\,.$ (11)
@@
    $E_{D}({\boldsymbol{\theta}}_{t})-E_{S}({\boldsymbol{\theta}}_{t})\leq\sqrt{\frac{\beta\operatorname{ess\,sup}_{p_{0}}L_{S}({\boldsymbol{\theta}})+\ln\left(1/\delta\right)}{2N}}\,.$ (12)
@@
The proof is simple — by assumption, in both cases $p_{0}=\nu$ so $D_{\infty}\left(p_{0}\,\|\,\nu\right)=0$. The rest is a direct substitution into Theorem 2.7.
3.1 Interpreting Corollary 3.1
@@
Corollary 3.1 raises questions on the relevance of this setting, which we address below: (1) How large is $\mathbb{E}_{p_{0}}L_{S}\left({\boldsymbol{\theta}}\right)$ in practically relevant cases? (2) Can we attribute the generalization to the regularization (either with the $\ell_{2}$ regularization term, or the bounded domain)? (3) Can models successfully train in the presence of noise with a variance large enough to make the bounds non-vacuous?
@@
Magnitude of the initial loss. Commonly, the dependence on $\mathbb{E}_{p_{0}}L_{S}\left({\boldsymbol{\theta}}\right)$ with realistic $p_{0}$ and $L_{S}$ is relatively mild. For example, using standard initialization schemes, Gaussian process approximations [41, 34, 30, 22] imply that the output of an infinitely wide fully connected neural network converges to a Gaussian with mean 0 and $O(1)$ variance at initialization. So in many cases $\mathbb{E}_{p_{0}}L_{S}\left({\boldsymbol{\theta}}\right)=O(1)$, such as for the scalar square and logistic losses. In the multi-output case, $\mathbb{E}_{p_{0}}L_{S}\left({\boldsymbol{\theta}}\right)$ may also depend on the number of outputs (e.g., logarithmically so in softmax-cross-entropy).
Magnitude of regularization. In the above result we must use regularization (or a bounded domain) that matches the initialization $p_{0}$ (this can be somewhat relaxed, see Section 3.2). The same assumption, that the regularization matches the initialization, was also made in other theoretical works on CLD [40, 32, 16]. Note that, NN models regularized this way remain highly expressive, both empirically (Appendix F) and theoretically (Appendix G), and therefore we cannot use this regularization alone, together with classical uniform convergence approaches to show generalization. Intuitively, this is because the regularization term can be tiny, for example, in (10) the regularization term is divided by $\beta$. Therefore, when $\beta=O\left(N\right)$ (which is sufficient for a non-vacuous result), $p_{0}=\nu$, and we use a standard deep nets initialization distribution $p_{0}$ (e.g., [18, 25], where $\lambda\propto{\mathrm{layer\,width}}$), the regularization coefficient is $O\left(\frac{\lambda}{N}\right)$ that is rather small in realistic cases. Therefore, we found (empirically) that it does not seem to have a significant effect at practical timescales.
@@
Magnitude of noise: theoretical perspective. In the above result we must use $\beta=O\left(N\right)$ to obtain a non-vacuous bound. This requirement is standard in many theoretical works. For example, as we will discuss below in Section 4.1, all previous generalization bounds for CLD and SGLD also required, to generalize well, $\beta=O\left(N\right)$ and potentially much worse (lower $\beta$). In addition, other theoretical works on noisy training also typically had $\beta=O\left(N\right)$ or worse. For example, when considering the ability of noisy gradient descent to escape saddle points, Jin et al. [27] uses noise sampled uniformly from a ball with a radius that depends on the dimensionality and smoothness of the problem, and thus cannot decay with $N$. Moreover, it is known that the Gibbs posterior [^3] Generalization bounds for the Gibbs posterior typically assume that it is “trained” and “tested” on the same function, while here the distribution is defined by the loss and “tested” on the error. generalizes well with $\beta=O\left(\sqrt{N}\right)$ (e.g. see Theorem 2.8 in 1), which is significantly smaller than $\beta=O\left(N\right)$. Lastly, in Appendix E we examine the impact of $\beta$ in the simple model of linear regression with i.i.d. standard Gaussian input, labels produced by a constant-magnitude teacher label noise, trained using regularized CLD as in (10), with $\lambda\propto d$ to match standard initialization. We find there that whenever $d\ll\beta\ll N$, the added noise does not significantly affect the training or population losses, and our bound is useful, i.e., it implies a vanishing generalization gap (since $\beta\ll N$ and $\mathbb{E}_{p_{0}}L=O(1)$). Note that $d\ll N$ is not a major constraint, since $d\ll N$ is required to obtain low population loss in this setting, even if we did not add noise to the training process (i.e. $\beta=\infty$).
@@
Magnitude of noise: empirical perspective. An inverse temperature of $\beta=O\left(N\right)$ is also relevant in many practical settings. For example, in Bayesian settings, when we wish to (approximately) sample from the posterior, it is quite common to use variants of SGLD; then inverse temperatures of order $\beta=O\left(N\right)$ are commonly used to achieve good generalization [52], which matches our results. In the standard practical training settings, the inverse temperature is a hyperparameter tuned to best fit a given problem. Empirically, in Appendix F we find that $\beta=O\left(N\right)$ can be tuned to obtain non-vacuous generalization bounds for overparameterized NNs in a few small binary classification datasets (binary MNIST, Fashion MNIST, SVHN, and a parity problem), i.e. the sum of the generalization gap bound and the training error is smaller than $0.5$. Importantly, these non-vacuous bounds do not use any trajectory-dependent quantities as other non-vacuous bounds (e.g. [13, 33]), which can make them arguably more useful as they can be calculated before training. The bounds are still not very tight (at noise levels that allow for non-vacuous bounds), but we believe there is still much room for improvement in future work.
@@ 3.2 Extensions and Modifications
where $\sigma^{2}\in\mathcal{C}^{2}$. For example, in Section D.1 we derive the explicit form of stationary distributions when $\sigma^{2}\left({\boldsymbol{\theta}}\right)=\left(L_{S}\left({\boldsymbol{\theta}}\right)+\alpha\right)^{k}$ or $\sigma^{2}\left({\boldsymbol{\theta}}\right)=e^{\alpha L\left(\alpha\right)}$, for some $k\in\mathbb{N}$ and $\alpha>0$. In both cases, the analytic form of the stationary potential $\Psi$ can be used directly with Theorem 2.7 to derive generalization bounds.
Restricted initialization. In Section D.2 we present generalizations of Corollary 3.1 to cases where $p_{0}$ and $\nu$ are different. Specifically, for the bounded case we consider $p_{0}$ that is uniform in a subset $\Theta_{0}\subset\Theta$ of the domain, and for the regularized case we consider general diagonal Gaussian initialization and regularization. In particular, this means that some of the parameters can be more loosely regularized/bounded at a cost proportional to their number. For example, in a deep NN, if only a single layer is loosely regularized/bounded, the KL-divergence cost will be proportional only to the number of parameters in that layer, not the entire $d$.
@@ 4 Related Work
A common type of generalization bounds consists of a measure of the dependence between the learned model and the dataset used to train it, such as the mutual information between the data and algorithm [45, 53, 48] or the KL-divergence between the predictor’s distribution and any data-independent distribution [36, 9, 1]. In particular, recent works were able to estimate such dependence measures from trained models to derive non-vacuous generalization bounds, even for deep overparameterized models. For example, Dziugaite et al. [14] used held-out data to bound the KL-divergence in a PAC-Bayes bound with a data dependent prior. Other works used some property of the trained model to estimate the information content, adding valuable insight to the mechanisms facilitating the successful generalization, such as the size of the compressed model after training, due to noise stability [3], and data structure [33].
@@ 4.1 Explicit Comparison for CLD
  • Theorem 2.7 and Corollary 3.1 demonstrate that our results hold for general initialization-regularization pairs, beyond Gaussian initialization with matching $\ell^{2}$ regularization.
In Table 1 we compare in more detail Corollary 3.1 to other bounds that remain bounded as $t\to\infty$.
@@ 4.2 Technical Novelty
As a representative example, we first focus on Raginsky et al. [46], which provided a bound for CLD (as an intermediate step for deriving a generalization bound for SGLD, a discretized version of CLD). Using spectral methods [e.g. 5], they bound the distance between the process’ distribution to the Gibbs posterior, which, when combined with the generalization bound for the Gibbs distribution, results in generalization bounds for intermediate times. Our Corollary 2.5 and the preceding arguments are similar to the proof of Lemma 3.4 of Raginsky et al. [46] that bounds the divergence between the initialization and the Gibbs distribution, where their dissipativity coefficient $m$ corresponds to our explicit $\ell^{2}$ regularization coefficient $\lambda$. We use some significant observations that make the bound simpler, and time/dimension/Lipschitz/smoothness independent.
@@ 4.3 Generalization Guarantees Applicable for Neural Networks
Many additional lines of work established generalization guarantees applicable for NNs, but are less directly related to our work. These results have some limitations that do not exist in ours. For example, NTK analysis [26] can imply generalization guarantees in certain settings, but they do not allow for feature learning; Mean-field results [37] require non-standard initialization and specific architectures; Algorithmic stability analysis [7, 23, 47, 31, 51] only apply when the number of iterations is sufficiently small; Norm-based generalization bounds [6, 19] ignore optimization aspects and depend exponentially on the network’s depth; And bounds for random interpolators [8] involve impractical training procedures.
@@ 5 Discussion, Limitations, and Future Work
where ${\boldsymbol{\Sigma}}$ is a matrix-valued dispersion coefficient. In contrast, in this paper, to derive concrete generalization bounds, we focused on CLD with isotropic noise, i.e. such that ${\boldsymbol{\Sigma}}$ is a scalar multiple of the identity matrix. The reason for this was that our bound (Corollary 3.1) relies on explicit analytical expressions or bounds on stationary distributions, which are difficult to find in the general case. In addition, in typical overparameterized settings, the noise induced by the randomness of SGD may not only be non-isotropic, but also low-rank. The analysis of such processes poses various challenges beyond the ability to derive an analytic form for their stationary distribution. For example, they may concentrate on low-dimensional manifolds, possibly making the KL-divergence term infinite, or making some of the assumptions unrealistic (e.g. the choice of initial distribution).
@@
No regularization. In this work, we only considered processes that have stationary probability measures. For this reason, in the examples in Section 3 we used either a bounded domain or regularization. This seems essential for generalization at $t\to\infty$, unless there are other architectural constrains. For example, consider training a model for classification of randomly labeled data. Without regularization, sufficiently expressive models are likely to arrive (at some point) at high training accuracy, yet it cannot generalize in this setting. Nonetheless, it might be possible to ensure generalization as a function of time, but here we focus on time-independent bounds.
@@
Discrete time steps. The behavior of SGD with a large step size may be qualitatively different than that of the continuous process considered here. Specifically, Azizian et al. [4] showed that while the asymptotic distribution of SGD resembles the Gibbs posterior, it is influenced by the step size, and geometry of the loss surface. While an extension of our analysis to this setting is straightforward given a stationary distribution, such stationary distributions are typically hard to find explicitly (except in simple cases, such as quadratic potentials), and the error terms coming from their approximations are typically detrimental to finding non-vacuous generalization bounds, as they may depend on the dimension of the parameters through the model’s Lipschitz or smoothness coefficients, etc. (40, 32, 16, 12). Hence, a direct application of our approach to such algorithms requires additional considerations. An alternative approach is to incorporate a Metropolis-Hastings type rejection [39, 24], ensuring that the stationary distribution is indeed the Gibbs posterior.
@@
Can noise be useful for generalization? There is a long line of work in the literature (e.g. see [17] and references therein), debating the effect of noise on generalization. Our work does not imply that higher noise improves the test error, only that it decreases the gap between training and testing. Since higher noise could hurt the training error, the overall effect depends on the specific situation. Even if introducing noise does not improve test performance, there could still be an advantage to introducing noise, based on our results, in that it reduces the gap and thus could increase the training error to match the test error in cases we cannot hope to learn (i.e. to get small test error). This is a good thing since it prevents being mislead by overfitting, hopefully without hurting the test error when we can generalize well (i.e. in learnable regimes, both training and test errors are low, perhaps also without noise, but in non-learnable regimes, where the test error is necessarily high, noise forces the training error to be high as well, so that the gap is small). Indeed, in our small-scale experiments in Appendix F, we noticed that a small amount of noise can decrease the generalization gap, without significantly harming the test error (e.g. see the bottom half of Tables 2, 3 and 4). Further analysis is necessary in order to establish general conditions under which test performance is not significantly hurt by noise, while ensuring a small gap. This, in particular, requires studying the effect of noise on the training loss, and what noise level still ensures obtaining small training loss in learnable regimes.

### v2

@@ 3 Special Case: Continuous Langevin Dynamics
Clearly, given Theorem˜2.7 all we need to do in order to derive explicit generalization bounds for any Markovian training procedure, is to find a stationary distribution, and bound its potential (or its expectation at $p_{0}$). In this section, we will exemplify our results in a few special cases of continuous-time Langevin dynamics (CLD), a commonly studied approximation for NN training with “infinitesimal learning rate” (e.g. [41], see Section˜4.1 for additional references), which have a normalized stationary distribution that we can write analytically.
@@
where $L_{S}\geq 0$ is twice continuously differentiable, $\mathbf{w}_{t}$ is a standard Brownian motion, and $\mathbf{r}_{t}$ is a reflection process that constrains ${\boldsymbol{\theta}}_{t}$ within $\Theta$. Such weight clipping is quite common in practical scenarios such as NN training. For simplicity, we assume that $\mathbf{r}_{t}$ has normal reflection, meaning that the reflection is perpendicular to the boundary. An established result in the analysis of SDERs states that under these assumptions (9) has a stationary distribution $p_{\infty}\left({\boldsymbol{\theta}}\right)\propto e^{-\beta L_{S}\left({\boldsymbol{\theta}}\right)}\mathbb{I}_{\Theta}\left\{{\boldsymbol{\theta}}\right\}$ (see Section˜H.2). Thus, when $p_{0}=\mathrm{Uniform}\left(\Theta\right)$, we have $p_{0}=\nu$.
@@
    $\mathbb{E}_{{\boldsymbol{\theta}}_{t}\sim p_{t}}\left[E_{D}({\boldsymbol{\theta}}_{t})-E_{S}({\boldsymbol{\theta}}_{t})\mid S\right]\leq\sqrt{\frac{\beta\mathbb{E}_{{\boldsymbol{\theta}}\sim p_{0}}\left[L_{S}({\boldsymbol{\theta}})\mid S\right]+\ln\left(N/\delta\right)}{2N}}\,.$ (11)
@@
    $E_{D}({\boldsymbol{\theta}}_{t})-E_{S}({\boldsymbol{\theta}}_{t})\leq\sqrt{\frac{\beta\operatorname{ess\,sup}_{p_{0}}L_{S}({\boldsymbol{\theta}})+\ln\left(N/\delta\right)}{2N}}\,.$ (12)
@@
The proof is simple — by assumption, in both cases $p_{0}=\nu$ so $D_{\infty}\left(p_{0}\,\|\,\nu\right)=0$. The rest is a direct substitution into Theorem˜2.7, and in particular, using $\beta L_{S}$ as potential $\Psi_{S}$.
3.1 Interpreting Corollary˜3.1
@@
Corollary˜3.1 raises questions on the relevance of this setting, which we address below: (1) How large is $\mathbb{E}_{p_{0}}L_{S}\left({\boldsymbol{\theta}}\right)$ in practically relevant cases? (2) Can we attribute the generalization to the regularization (either with the $\ell_{2}$ regularization term, or the bounded domain)? (3) Can models successfully train in the presence of noise with a variance large enough to make the bounds non-vacuous?
@@
Magnitude of the initial loss. Commonly, the dependence on $\mathbb{E}_{p_{0}}L_{S}\left({\boldsymbol{\theta}}\right)$ with realistic $p_{0}$ and $L_{S}$ is relatively mild. For example, using standard initialization schemes, Gaussian process approximations [50, 42, 35, 25] imply that the output of an infinitely wide fully connected neural network converges to a Gaussian with mean 0 and $O(1)$ variance at initialization. So in many cases $\mathbb{E}_{p_{0}}L_{S}\left({\boldsymbol{\theta}}\right)=O(1)$, such as for the scalar square and logistic losses. In the multi-output case, $\mathbb{E}_{p_{0}}L_{S}\left({\boldsymbol{\theta}}\right)$ may also depend on the number of outputs (e.g., logarithmically so in softmax-cross-entropy). A more difficult question is concerned with the case that $\operatorname{ess\,sup}_{p_{0}}L_{S}=\infty$, which is common when $p_{0}$ has infinite support. This can be mitigated by clipping the loss, which is standard in practice (e.g. in reinforcement learning [48, 62]) and in the theory of optimization [37, 33]. Moreover, this clipping can be done in a differentiable way (e.g. using either softmin, tanh (e.g. $c\cdot\tanh(L/c)$), etc) and at values only slightly higher than the typical loss at the initialization (since the loss is roughly monotonically decreasing in CLD with small noise, the optimization process would typically operate below the clipping and will not be affected by it).
[…]
@@
Magnitude of noise: theoretical perspective. In the above result we must use $\beta=O\left(N\right)$ to obtain a non-vacuous bound. This requirement is standard in many theoretical works. For example, as we will discuss below in Section˜4.1, all previous generalization bounds for CLD and SGLD also required, to generalize well, $\beta=O\left(N\right)$ and potentially much worse (lower $\beta$). In addition, other theoretical works on noisy training also typically had $\beta=O\left(N\right)$ or worse. For example, when considering the ability of noisy gradient descent to escape saddle points, Jin et al. [30] uses noise sampled uniformly from a ball with a radius that depends on the dimensionality and smoothness of the problem, and thus cannot decay with $N$. Moreover, it is known that the Gibbs posterior [^7] Generalization bounds for the Gibbs posterior typically assume that it is “trained” and “tested” on the same function, while here the distribution is defined by the loss and “tested” on the error. generalizes well with $\beta=O\left(\sqrt{N}\right)$ (e.g. see Theorem 2.8 in 1), which is significantly smaller than $\beta=O\left(N\right)$. Lastly, in Appendix˜E we examine the impact of $\beta$ in the simple model of linear regression with i.i.d. standard Gaussian input, labels produced by a constant-magnitude teacher label noise, trained using regularized CLD as in (10), with $\lambda\propto d$ to match standard initialization. We find there that whenever $d\ll\beta\ll N$, the added noise does not significantly affect the training or population losses, and our bound is useful, i.e., it implies a vanishing generalization gap (since $\beta\ll N$ and $\mathbb{E}_{p_{0}}L=O(1)$). Note that $d\ll N$ is not a major constraint, since $d\ll N$ is required to obtain low population loss in this setting, even if we did not add noise to the training process (i.e. $\beta=\infty$).
@@
Magnitude of noise: empirical perspective. An inverse temperature of $\beta=O\left(N\right)$ is also relevant in many practical settings. For example, in Bayesian settings, when we wish to (approximately) sample from the posterior, it is quite common to use variants of SGLD; then inverse temperatures of order $\beta=O\left(N\right)$ are commonly used to achieve good generalization [69], which matches our results. In the standard practical training settings, the inverse temperature is a hyperparameter tuned to best fit a given problem. Empirically, in Appendix˜F we find that $\beta=O\left(N\right)$ can be tuned to obtain non-vacuous generalization bounds for overparameterized NNs in a few small binary classification datasets (binary MNIST, Fashion MNIST, SVHN, and a parity problem), i.e. the sum of the generalization gap bound and the training error is smaller than $0.5$. Importantly, these non-vacuous bounds do not use any trajectory-dependent quantities as other non-vacuous bounds (e.g. [15, 39]), which can make them arguably more useful as they can be calculated before training. The bounds are still not very tight (at noise levels that allow for non-vacuous bounds), but we believe there is still much room for improvement in future work.
@@ 3.2 Extensions and Modifications
where $\sigma^{2}\in\mathcal{C}^{2}$. For example, in Section˜D.1 we derive the explicit form of stationary distributions when $\sigma^{2}\left({\boldsymbol{\theta}}\right)=\left(L_{S}\left({\boldsymbol{\theta}}\right)+\alpha\right)^{k}$ or $\sigma^{2}\left({\boldsymbol{\theta}}\right)=e^{\alpha L_{S}\left({\boldsymbol{\theta}}\right)}$, for some $k\in\mathbb{N}$ and $\alpha>0$. In both cases, the analytic form of the stationary potential $\Psi$ can be used directly with Theorem˜2.7 to derive generalization bounds.
Restricted initialization. In Section˜D.2 we present generalizations of Corollary˜3.1 to cases where $p_{0}$ and $\nu$ are different. Specifically, for the bounded case we consider $p_{0}$ that is uniform in a subset $\Theta_{0}\subset\Theta$ of the domain, and for the regularized case we consider general diagonal Gaussian initialization and regularization. In particular, this means that some of the parameters can be more loosely regularized/bounded at a cost proportional to their number. For example, in a deep NN, if only a single layer is loosely regularized/bounded, the KL-divergence cost will be proportional only to the number of parameters in that layer, not the entire $d$.
@@ 4 Related Work
A common type of generalization bounds consists of a measure of the dependence between the learned model and the dataset used to train it, such as the mutual information between the data and algorithm [58, 70, 61] or the KL-divergence between the predictor’s distribution and any data-independent distribution [44, 9, 1]. In particular, recent works were able to estimate such dependence measures from trained models to derive non-vacuous generalization bounds, even for deep overparameterized models. For example, Dziugaite et al. [17] used held-out data to bound the KL-divergence in a PAC-Bayes bound with a data-dependent prior. Other works used some property of the trained model to estimate the information content, adding valuable insight to the mechanisms facilitating the successful generalization, such as the size of the compressed model after training, due to noise stability [3], and data structure [39].
@@ 4.1 Explicit Comparison for CLD
  • Theorem˜2.7 and Corollary˜3.1 demonstrate that our results hold for general initialization-regularization pairs, beyond Gaussian initialization with matching $\ell^{2}$ regularization.
In Table˜1 we compare in more detail Corollary˜3.1 to other bounds that remain bounded as $t\to\infty$.
@@ 4.2 Technical Novelty
As a representative example, we first focus on Raginsky et al. [59], which provided a bound for CLD (as an intermediate step for deriving a generalization bound for SGLD, a discretized version of CLD). Using spectral methods [e.g. 5], they bound the distance between the process’ distribution to the Gibbs posterior, which, when combined with the generalization bound for the Gibbs distribution, results in generalization bounds for intermediate times. Our Corollary˜2.5 and the preceding arguments are similar to the proof of Lemma 3.4 of Raginsky et al. [59] that bounds the divergence between the initialization and the Gibbs distribution, where their dissipativity coefficient $m$ corresponds to our explicit $\ell^{2}$ regularization coefficient $\lambda$. We use some significant observations that make the bound simpler, and time/dimension/Lipschitz/smoothness independent.
@@ 4.3 Generalization Guarantees Applicable for Neural Networks
Many additional lines of work established generalization guarantees applicable for NNs, but are less directly related to our work. These results have some limitations that do not exist in ours. For example, NTK analysis [29] can imply generalization guarantees in certain settings, but they do not allow for feature learning; Mean-field results [45] require non-standard initialization and specific architectures; Algorithmic stability analysis Bousquet and Elisseeff [7], Hardt et al. [26], Richards and Rabbat [60], Lei et al. [36], Wang et al. [67] only apply when the number of iterations is sufficiently small; Norm-based generalization bounds [6, 22] ignore optimization aspects and depend exponentially on the network’s depth; And bounds for random interpolators [8] involve impractical training procedures.
A closely related setting to the one studied here is SGLD, i.e. a discretized version of CLD. There is an extensive line of work bounding the generalization gap of such models (see [59, 49, 55, 51, 18, 19, 13] for a partial list). These results typically have a significant dependence on hyperparameter stemming from the discretization such as the learning rate and batch size, or suffer from constraints similar to the ones discussed in Section˜4.1, such as dependence on trajectory or dimensionality (e.g. via smoothness, parameter norms, log-Sobolev or spectral gap constants).
@@ 5 Discussion, Limitations, and Future Work
where $\boldsymbol{\Sigma}$} is a matrix-valued dispersion coefficient. In contrast, in this paper, to derive concrete generalization bounds, we focused on CLD with isotropic noise, i.e. such that $\boldsymbol{\Sigma}$} is a scalar multiple of the identity matrix. The reason for this was that our bound (Corollary˜3.1) relies on explicit analytical expressions or bounds on stationary distributions, which are difficult to find in the general case. In addition, in typical overparameterized settings, the noise induced by the randomness of SGD may not only be non-isotropic, but also low-rank. The analysis of such processes poses various challenges beyond the ability to derive an analytic form for their stationary distribution. For example, they may concentrate on low-dimensional manifolds, possibly making the KL-divergence term infinite, or making some of the assumptions unrealistic (e.g. the choice of initial distribution).
@@
No regularization. In this work, we only considered processes that have stationary probability measures. For this reason, in the examples in Section˜3 we used either a bounded domain or regularization. This seems essential for generalization at $t\to\infty$, unless there are other architectural constrains. For example, consider training a model for classification of randomly labeled data. Without regularization, sufficiently expressive models are likely to arrive (at some point) at high training accuracy, yet it cannot generalize in this setting. Nonetheless, it might be possible to ensure generalization as a function of time, but here we focus on time-independent bounds.
@@
Discrete time steps. The behavior of SGD with a large step size may be qualitatively different than that of the continuous process considered here. Specifically, Azizian et al. [4] showed that while the asymptotic distribution of SGD resembles the Gibbs posterior, it is influenced by the step size, and geometry of the loss surface. While an extension of our analysis to this setting is straightforward given a stationary distribution, such stationary distributions are typically hard to find explicitly (except in simple cases, such as quadratic potentials), and the error terms coming from their approximations are typically detrimental to finding non-vacuous generalization bounds, as they may depend on the dimension of the parameters through the model’s Lipschitz or smoothness coefficients, etc. (49, 38, 19, 14). Hence, a direct application of our approach to such algorithms requires additional considerations. An alternative approach is to incorporate a Metropolis-Hastings type rejection [47, 27], ensuring that the stationary distribution is indeed the Gibbs posterior.
@@
Can noise be useful for generalization? There is a long line of work in the literature (e.g. see [20] and references therein), debating the effect of noise on generalization. Our work does not imply that higher noise improves the test error, only that it decreases the gap between training and testing. Since higher noise could hurt the training error, the overall effect depends on the specific situation. Even if introducing noise does not improve test performance, there could still be an advantage to introducing noise, based on our results, in that it reduces the gap and thus could increase the training error to match the test error in cases we cannot hope to learn (i.e. to get small test error). This is a good thing since it prevents being mislead by overfitting, hopefully without hurting the test error when we can generalize well (i.e. in learnable regimes, both training and test errors are low, perhaps also without noise, but in non-learnable regimes, where the test error is necessarily high, noise forces the training error to be high as well, so that the gap is small). Indeed, in our small-scale experiments in Appendix˜F, we noticed that a small amount of noise can decrease the generalization gap, without significantly harming the test error (e.g. see the bottom half of Tables˜2, 3 and 4). Further analysis is necessary in order to establish general conditions under which test performance is not significantly hurt by noise, while ensuring a small gap. This, in particular, requires studying the effect of noise on the training loss, and what noise level still ensures obtaining a small training loss in learnable regimes.
Acknowledgments and Disclosure of Funding
The research of DS was Funded by the European Union (ERC, A-B-C-Deep, 101039436). Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency (ERCEA). Neither the European Union nor the granting authority can be held responsible for them. DS also acknowledges the support of the Schmidt Career Advancement Chair in AI. GV is supported by the Israel Science Foundation (grant No. 2574/25), by a research grant from Mortimer Zuckerman (the Zuckerman STEM Leadership Program), and by research grants from the Center for New Scientists at the Weizmann Institute of Science, and the Shimon and Golde Picker – Weizmann Annual Grant. Part of this work was done as part of the NSF-Simons funded Collaboration on the Mathematics of Deep Learning. NS was partially supported by the NSF TRIPOD Institute on Data Economics Algorithms and Learning (IDEAL) and an NSF-IIS award.
Appendix structure:
  • In Appendix˜A we recap and establish notation and conventions, and present some well-known lemmas.
  • In Appendix˜B we prove Theorem˜2.7 and its related claims in Section˜2.
  • In Appendix˜C we discuss the tightness and necessity of the divergence conditions found in Appendix˜B.
  • In Appendix˜D we prove a generalized version of Corollary˜3.1.
  • The bounds found in this paper only bound the generalization gap, and not the absolute error of a model. In Appendix˜E and Appendix˜F we study the applicability of our bound in realistic settings. Specifically, whether the regime in which the bound on the generalization gap is non-vacuous allows for meaningful learning, i.e. coincides with a regime in which the absolute training error is also small. In Appendix˜E we study linear regression trained with CLD, for which we can analytically characterize the training loss, and in Appendix˜F we experiment with NNs trained with SGLD (discretized version of CLD) on standard training sets.
  • As Section˜3 deals only with models trained with some form of regularization, it is natural to ask whether the regularization alone is sufficient for the use of uniform convergence to arrive the desired generalization bounds. In Appendix˜G we show that the regularization used in Section˜3 is not sufficient for such bounds, and that the models can remain highly expressive.
  • Finally, for completeness, in Appendix˜H we recall some definitions and properties related to SDEs used throughout the paper.